(best viewed in Firefox)
Add Yourself to the Atacama Mailing List
Salar Grande, Atacama Desert, Chile
- Science operations during traverse from Pampa Nevada site to next site
- Deploy spectrometer and fluorescence microscope
Status and Progress
- Experienced narrow bandwidth. The day started and ended difficulties
due to the narrow bandwidth of communication to the remote science
team. We have been experimenting with various communication systems
for connectivity to Hyperion (using VHF, 900MHz, and 2.4GHz 802.11b)
but have not address direct to satellite communication. As a result we
have been up/downloading with Hyperion and then transferring data
(including these reports) over a 56k internet link. It works fine but
this morning we had our incoming pipe jam up and this evening it was
down entirely. This has necessitated two trips into Iquique.
Nevertheless we completed the full science plan generated by the remote
team at NASA Ames.
- Improved compression. We worked out a way to further compress many
of the images we are taking for panoramic imagery and still allow
reconstruction in Eventscope. In some cases this has resulted in a 15x
reduction in data size. We are also working on more efficient ways to
specify rover commands to improve operations.
- Conducted remote science investigation. Today's science plan called
for high-resolution images, spectra, and flourescence microscopic
images of eight targets or general locations. Hyperion also traversed
750m from the Pampa Nevada "landing site" to the base of an scarp,
taking periodic panoramic image datasets. We have been very interested
in the south side of Pampa Nevada where Hyperion was operating. This
area is shaded in the morning and appears to have higher moisture
content/lower insolation. There is lichen present on some slopes.
With Hyperion on traverse we moved in with the manually deployed
spectrometer and microscope (to be integrated next year) and for
ground-truth sampling. We have found material directly on the salar
that appears moss-like in texture (see images in the gallery) in colors
of green, red/brown and yellow/brown. However spectra do not show
strong organic content and fluorescence microscopy reveals minimal
fluorescence. Further data collection and investigation is needed.
The remote science team is analyzing data independently while charting
a course southwest to an alluvial fan that they have identified in
satellite imagery. Remote science operations continue through Friday.
- Science operations at Beacon site
- Collect localization data from Sun sensor and side-view camera
- Test mission planning of science instrument (SPI) data collection
Morning: Clear, light dew
Evening: High clouds, nice sunset, breezy
Weather data is from April 21, the last complete day of data.
Vulture Count: 5
Quote of the Day
for i in `ls -l | cut -d ' ' -f 20`; do head $i -n 2 | tail -n 1 | cut
-d '#' -f 2; done > sol2-sample10.txt